# SenseShift6D: Multimodal RGB-D Benchmarking for Robust 6D Pose Estimation across Environment and Sensor Variations

### v1

@@ Abstract
Recent advances on 6D object-pose estimation has achieved high performance on representative benchmarks such as LM-O, YCB-V, and T-Less. However, these datasets were captured under fixed illumination and camera settings, leaving the impact of real-world variations in illumination, exposure, gain or depth-sensor mode – and the potential of test-time sensor control to mitigate such variations – largely unexplored. To bridge this gap, we introduce SenseShift6D, the first RGB-D dataset that physically sweeps 13 RGB exposures, 9 RGB gains, auto-exposure, 4 depth-capture modes, and 5 illumination levels. For three common household objects (spray, pringles, and tincase), we acquire 101.9k RGB and 10k depth images, which can provide 1,380 unique sensor-lighting permutations per object pose. Experiments with state-of-the-art models on our dataset show that applying sensor control during test-time induces greater performance improvement over digital data augmentation, achieving performance comparable to or better than costly increases in real-world training data quantity and diversity. Adapting either RGB or depth sensors individually is effective, while jointly adapting multimodal RGB-D configurations yields even greater improvements. SenseShift6D extends the 6D-pose evaluation paradigm from data-centered to sensor-aware robustness, laying a foundation for adaptive, self-tuning perception systems capable of operating robustly in uncertain real-world environments.
Our dataset is available at: huggingface.co/datasets/Yegyu/SenseShift6D
Associated scripts can be found at: github.com/yegyu-han/SenseShift6D
@@ 1 Introduction
Estimating the 6D pose of everyday objects is a critical computer vision task, enabling fine-grained interactions in mixed-reality headsets, autonomous-vehicle manipulation, and embodied AI agents navigating cluttered environments  [5, 28, 24, 26, 16, 21]. The field has made substantial progress, with state-of-the-art methods achieving high accuracy on established benchmarks [7, 3, 10, 15, 11].
However, these impressive results are achieved under strictly controlled acquisition settings. Real-world deployments, in contrast, face variations: ambient illumination changes throughout the day, camera exposure and gain settings are adjusted by on-board auto-control or user preference, and commodity depth sensors can also switch capture modes on purpose. These environmental and sensor dynamics introduce complex artifacts, including nonlinear color shifts and structured depth noise, which are challenging to replicate accurately with synthetic augmentations or compensate for during post-processing [1]. Consequently, existing benchmarks – collected with fixed illumination, exposure, gain, and depth configurations – cannot adequately represent real-world scenarios. Therefore, the research community lacks a systematic benchmark for quantifying how 6D-pose estimators respond to per-frame sensor variation or for evaluating emerging test-time sensor-control strategies [1, 2] designed to mitigate these shifts.
Recent attempts to tackle this robustness gap have explored adaptive sensor control, which dynamically adjusts sensor parameters at test time to maximize accuracy without modifying the underlying model. Studies in image classification have demonstrated that test-time sensor control improves model robustness more effectively than scaling up the model size or augmenting training data [1, 2]. However, these efforts have been limited to unimodal RGB tasks. Importantly, no prior work has examined multimodal (RGB and depth) sensor control or provided a benchmark that physical sweeps multimodal sensor parameters with ground-truth 6D pose. Consequently, the effectiveness of adaptive multimodal sensing and its interplay with traditional data-centric training remain an open question.
To bridge this gap, we introduce SenseShift6D, a novel RGB-D benchmark for 6D pose estimation that physically sweeps key sensor and lighting parameters. Using an Intel RealSense RGB-D camera in a controlled darkroom setup, we systematically vary 13 exposure levels, 9 gain settings, 4 depth-capture modes, and 5 illumination conditions across three representative household objects: a textured spray can, a cylindrical Pringles tube, and a glossy tin case. This protocol produces a dataset of 101.9k RGB and 10k depth frames, which can provide 1,380 unique sensor-lighting permutations per object pose (as illustrated in Figures 1 and 2), each annotated via 6D ground-truth via ArUco calibration and manual refinement (with $\sim$3 mm errors). For controlled study, experiments in this paper are conducted on the four subsets of the entire dataset: Train-Def, recorded under default camera and illumination settings, and Train-Var, covering a comprehensive multimodal parameter grid (8 exposure, 5 gain, auto-exposure, 4 depth-capture settings with 5 lighting conditions). Correspondingly, we construct two test splits (Test-Def and Test-Var) without overlap with the training configurations. As all variations are physically realized, the captured images preserve authentic sensor artifacts, including nonlinear color shifts, photon-shot noise, and structured depth errors, that synthetic augmentations cannot replicate. Thus, SenseShift6D establishes the first realistic testbed for sensor-aware 6D pose estimation and adaptive sensing.
@@
  • Robustness through test-time sensor control: We demonstrate that adjusting sensor parameters during inference under environmental variations improves 6D pose estimation accuracy by an average of 11.4% compared to standard auto-exposure with conventional digital augmentation. This indicates that sensor-level tuning can better account for the real-world variability faced during deployment. Moreover, sensor control proves to be a data-efficient approach to improving robustness, achieving comparable generalization with significantly less training data.
  • Potential of multimodal sensor adaptation: While adapting either RGB or depth sensors individually already yields noticeable improvements, jointly optimizing both RGB and depth parameters at test-time provides even greater gains. This emphasizes the value of studying multimodal sensor interactions and highlights new directions for advancing the field of 6D pose estimation.
@@ 2 Related Work
Dataset Environmental and sensor variations Objects RGB frames Depth frames
@@
LM (ACCV ’12) [7] $\times$ $\times$ $\times$ $\times$ 15 18.2k 18.2k
LM-O (ECCV ’14) [3] $\times$ $\times$ $\times$ $\times$ 8 1.2k 1.2k
T-LESS (WACV ’17) [10] $\times$ $\times$ $\times$ $\times$ 30 49k 49k
YCB-V (RSS ’18) [27] $\times$ $\times$ $\times$ $\times$ 21 133k 133k
TUD-L (ECCV ’18) [11] 8 locations $\times$ $\times$ $\times$ 3 62k 62k
HB (ICCVW ’19) [15] 2 locations $\times$ $\times$ $\times$ 33 34.8k 34.8k
HOPE (IROS ’22) [23] 5 locations $\times$ $\times$ $\times$ 28 238 238
IPD (CVPR ’24) [14] 3 locations 4 levels $\times$ $\times$ 20 30k 30k
SenseShift6D (Ours) 5 levels 13 levels 9 levels 4 modes 3 101.9k 10k
@@
Classic datasets. LineMOD (LM) [7] and its occlusion subset (LM-O) [3] pioneered RGB-D pose estimation benchmarks in controlled tabletop setups. Subsequent benchmarks such as T-LESS [10], which provided a large collection of texture-less industrial objects, and YCB-V [27], which offered cluttered scenes with temporal sequences, expanded the scope of evaluation. While these benchmarks provide robust evaluation regarding variations in objects and their poses, all were captured under fixed illumination and camera settings.
Initial exploration of environmental and sensor variation. TUD-L [11], HomebrewedDB (HB) [15], and HOPE [23] introduced variations in illumination conditions but retained fixed camera settings with auto-exposure enabled. The recent Industrial Parts Dataset (IPD) [14] was the first benchmark to vary camera parameters, capturing four exposure levels under multiple lighting conditions. However, IPD focuses on texture-less industrial parts and does not explore variations in RGB gain or depth-sensor modes, thus limiting its suitability for adaptive-sensing research.
@@ 2.2 Adaptive Sensor Control Benchmark
Recent studies have explored test-time adaptation of camera sensor parameters to improve downstream vision performance. ImageNet-ES [1] and ImageNet-ES-Diverse [2] captured real-world scenes using a physical camera, systematically varying aperture, shutter speed, and ISO across multiple lighting conditions for image classification tasks. These benchmarks demonstrated that physical variation cannot be replicated by synthetic augmentations and dynamically selecting optimal camera parameters at test time can significantly increase the performance of lightweight models, enabling accuracy competitive with larger foundation models without additional training.
@@ 3 SenseShift6D
Table 2: SenseShift6D Configurations for Data split.
Data split Brightness (%) Sensor configuration Exposure ($\mu$s) Gain Depth-capture mode Captured images
Train-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 385, Depth: 385 1 RGB-D scene/pose
Train-Var 5, 25, 50, 75, 100 Auto + Manual (41 RGB options, 1 depth options) Auto, 2, 4, 19, 78, 312, 1250, 5000, 10000 Auto, 0, 32, 64, 96, 128 Default RGB: 80.0k, Depth: 1.9k 205 RGB-D scenes/pose
Test-Def 50 Auto (1 RGB option, 1 depth option) Auto Auto Default RGB: 128, Depth: 128 1 RGB-D scenes/pose
Test-Var 5, 25, 50, 75, 100 Auto + Manual (21 RGB options, 4 Depth options) Auto, 9, 39, 156, 625, 2500 Auto, 16, 48, 80, 112 Default, High Accuracy, High Density, Medium Density RGB: 13k, Depth: 2.5k 420 RGB-D scenes/pose
Figure 3: Three household objects collected in SenseShift6D.
We introduce SenseShift6D, the first RGB-D benchmark that physically sweeps camera and lighting parameters, enabling rigorous evaluation of 6D object pose estimators under realistic environmental and sensor dynamics. Using an Intel RealSense D455 mounted in a darkroom, we systematically capture three representative household objects – a textured spray can, a cylindrical Pringles tube, and a glossy tin case (Figure 3) – under auto-exposure, 13 RGB-exposure times, 9 RGB-gain settings, 4 depth-capture presets, and 5 illumination levels (details in Table 2). Our dataset consists of 101.9k RGB and 10k depth frames, which can yield 1,380 unique sensor-lighting permutations per object pose. Accurate ground truth is obtained via a rigidly mounted ChArUco board and careful hand-eye calibration.
@@
  • Train-Var: Expands Train-Def, covering (8 exposures $\times$ 5 gains + 1 auto-exposure) $\times$ 1 depth modes $\times$ 5 brightness levels (205 RGB-D scenes/pose).
@@
  • Test-Var: Expands Test-Def, featuring (5 exposures $\times$ 4 gains + 1 auto-exposure) configurations, spanning the same depth-mode and illumination range (420 RGB-D scenes/pose).
As all variations are physically captured rather than digitally simulated, the resulting images inherently retain authentic artifacts, such as non-linear color shifts, photon-shot noise, and depth-mode-specific errors, that synthetic augmentation methods cannot reproduce [1]. SenseShift6D therefore offers the first realistic testbed for analyzing environment- and sensor-aware robustness, enabling side-by-side evaluation of per-frame adaptive-sensing policies against traditional data-centric training strategies.
@@ 3.1 Environmental and Sensor Variations
  • Exposure grid. The D455 supports exposure settings ranging from 1 $\mu$s to 10,000 $\mu$s. Empirical tests indicated that exposure times below 1,000 $\mu$s yield image quality comparable to the camera’s auto-exposure setting in well-lit environments. Given that this threshold is relatively low within the configurable range, we chose 13 logarithmically spaced exposure values to maximize coverage across practical operating conditions.
@@
Depth capture modes. The Intel RealSense D455 generates depth maps via stereo matching, estimating depth by aligning left and right camera images. This process relies on multiple internal parameters whose individual effects are complex and challenging to fine-tune directly. To manage this complexity, we leverage four predefined depth presets provided by Intel, each optimized for different usage scenarios [12]:
@@ 3.2 Dataset Collection
Table 3: GT annotation accuracy (mm).
@@
For automatic collection, the turntable was remotely controlled, and deliberate delays were introduced after changing lighting brightness or sensor parameters, allowing sufficient stabilization time. We employed one smart IR remote-control hub to send commands to the turntable and another hub to control three smart bulbs. Both hubs received commands from a cloud-based server, which coordinated the entire data-capture capture sequence through an integrated control script (see Figure 4).
@@ 3.3 Data Processing
Ground-Truth annotation. For each pose, we used the 50%-brightness, averaged auto-exposure RGB frame and went through the annotation process below, resulting in 2$\sim$3 mm errors.
  • Object masks: Initially, rough bounding boxes were generated using Grounding DINO [18], guided by shape-aware textual prompts tailored to each object. Boxes were manually reviewed and corrected if needed. The refined boxes were passed to FastSAM [29] to generate precise instance-level masks.
  • Object poses: An initial camera-to-object transformation was provided by detecting the ChArUco markerboard rigidly attached to the turntable. The object poses were then refined precisely in 3D using the annotation utility [6] in the BOP toolkit [9] with depth maps captured in High Density mode. Finally, all refined poses were re-projected onto the RGB images and visually inspected to confirm alignment accuracy.
  • Accuracy of GT annotation: Following the protocol in [10, 15], we calculate the difference $\delta=d_{c}-d_{r}$, where $d_{c}$ denotes the captured depth map and $d_{r}$ denotes the rendered depth map using the GT pose. Differences over 5 cm are treated as outliers and omitted. Table 4 reports the mean $\mu_{\delta}$, standard deviation $\sigma_{\delta}$, and the mean and median of the absolute differences $\mu_{|\delta|}$ and $med_{|\delta|}$. High-accuracy depth sensing is employed for reliable evaluation of GT accuracy analysis.

### v2

@@ Abstract
Recent advances on 6D object-pose estimation have achieved high performance on representative benchmarks such as LM-O, YCB-V, and T-Less. However, these datasets were captured under fixed illumination and camera settings, leaving the impact of real-world variations in illumination, exposure, gain or depth-sensor mode—and the potential of test-time sensor control to mitigate such variations—largely unexplored. To bridge this gap, we introduce SenseShift6D, the first RGB-D dataset that physically sweeps 13 RGB exposures, 9 RGB gains, auto-exposure, 4 depth-capture modes, and 5 illumination levels. For five common household objects (spray, pringles, tincase, sandwich, and mouse), we acquire 166.4k RGB and 16.7k depth images, which can provide 1,380 unique sensor-lighting permutations per object pose. Experiments with state-of-the-art models on our dataset demonstrate that applying multimodal sensor control at test time yields substantial performance gains, achieving a 19.5 pp improvement on pretrained generalizable models. It also enhances robustness precisely where those models tend to fail. Moreover, even instance-level pose estimators, where train and test set share identical object and background, performance still varies under environmental and sensor change, demonstrating that test-time sensor control remains effective compared to costly expansions in the quantity and diversity of real-world training data, without any additional training. SenseShift6D extends the object pose evaluation paradigm from data-centered to sensor-aware robustness, laying a foundation for adaptive, self-tuning perception systems capable of operating robustly in uncertain real-world environments.
@@ 1 Introduction
Estimating the 6D pose of everyday objects is a critical computer vision task, enabling fine-grained interactions in mixed-reality headsets, autonomous-vehicle manipulation, and embodied AI agents navigating cluttered environments  (Doughty & Ghugre, 2022; Zhang et al., 2023; Trojak et al., 2025; Wu et al., 2019; Ke et al., 2020; Ponimatkin et al., 2025). The field has made substantial progress, with state-of-the-art methods achieving high accuracy on established benchmarks (Hinterstoisser et al., 2012a; Brachmann et al., 2014; Hodan et al., 2017; Kaskman et al., 2019; Hodan et al., 2018b).
However, these impressive results are achieved under strictly controlled acquisition settings. Real-world deployments, in contrast, face substantial variations: ambient illumination changes throughout the day; camera exposure and gain are adjusted by on-board auto-control or user preference; and commodity depth sensors switch capture modes on purpose. These environmental and sensor dynamics introduce complex artifacts—nonlinear color shifts and structured depth noise—that are difficult to reproduce via synthetic augmentations or compensate for during post-processing (Baek et al., 2024). Consequently, existing benchmarks—mostly collected with fixed illumination, exposure, gain, and depth configurations—cannot adequately represent real-world scenarios. Therefore, the research community lacks a systematic benchmark to quantify how 6D-pose estimators respond to per-frame sensor variation or to evaluate emerging test-time sensor-control strategies (Baek et al., 2024; 2025) designed to mitigate these shifts.
Recent attempts to tackle this robustness gap have explored adaptive sensor control, which dynamically adjusts sensor parameters at test time to maximize accuracy without modifying the underlying model. Studies in image classification have demonstrated that test-time sensor control improves model robustness more effectively than scaling up the model size or augmenting training data (Baek et al., 2024; 2025). However, these efforts have been limited to unimodal RGB tasks. Importantly, no prior work has examined multimodal (RGB and depth) sensor control or provided a benchmark that physically sweeps multimodal sensor parameters with ground-truth 6D pose. Consequently, the effectiveness of adaptive multimodal sensing—and its interplay with traditional data-centric training—remains an open question.
@@
To bridge this gap, we introduce SenseShift6D, a novel RGB-D benchmark for 6D pose estimation that physically sweeps key sensor and lighting parameters. Using an Intel RealSense RGB-D camera in a controlled darkroom setup, we systematically vary 13 exposure levels, 9 gain settings, 4 depth-capture modes, and 5 illumination conditions across five representative household objects: a spray bottle, a Pringles tube, a tincase, a sandwich replica, and a vertical mouse. This protocol produces a dataset of 166.4k RGB and 16.7k depth frames, which can provide 1,380 unique sensor-lighting permutations per object pose (Figures 1 and 2), and 665.9k paired RGB-D scenes—to our knowledge, the largest of its kind. Each scene is annotated via 6D ground-truth via ArUco calibration and manual refinement (with $\sim$2 mm errors).
For controlled study, experiments in this paper are conducted on the four subsets of the entire dataset: Train-Def, captured under default camera and illumination settings, and Train-Var, covering a comprehensive multimodal parameter grid (8 exposure, 5 gain, auto-exposure, 1 depth-capture settings with 5 lighting conditions). Correspondingly, we construct two test splits (Test-Def and Test-Var) without overlap with the training configurations. As all variations are physically realized, the captured images preserve authentic sensor artifacts, including nonlinear color shifts, photon-shot noise, and structured depth errors, that synthetic augmentations cannot replicate. Thus, SenseShift6D establishes the first realistic testbed for sensor-aware 6D pose estimation and adaptive sensing.
@@
  • New Axes of Generalization: We reveal that environmental illumination and sensor settings constitute an orthogonal axis of generalization that current approaches do not address: (i) pretrained, generalizable models targeting unseen objects, occlusion, and clutter still degrade markedly under these shifts; and (2) instance-level models trained with standard augmentations likewise degrade even for known objects.
  • Robustness via multimodal test-time sensor control: We demonstrate that test-time sensor control substantially improves performance without model retraining. While single-modal control is already effective (e.g., +12.3 percentage points (pp) on pretrained models), multimodal control delivers further gains (e.g., +19.5 pp on the same models), highlighting the synergy from jointly adapting RGB and depth sensors.
@@ 2 Related Work
Dataset Environmental and sensor variations Objects RGB Depth RGB-D
@@
LM (Hinterstoisser et al., 2012a) $\times$ $\times$ $\times$ $\times$ 15 18.2k 18.2k 18.2k
LM-O (Brachmann et al., 2014) $\times$ $\times$ $\times$ $\times$ 8 1.2k 1.2k 1.2k
T-LESS (Hodan et al., 2017) $\times$ $\times$ $\times$ $\times$ 30 49k 49k 49k
YCB-V (Xiang et al., 2017) $\times$ $\times$ $\times$ $\times$ 21 133k 133k 133k
TUD-L (Hodan et al., 2018b) 8 options $\times$ $\times$ $\times$ 3 62k 62k 62k
HB (Kaskman et al., 2019) 2 options $\times$ $\times$ $\times$ 33 34.8k 34.8k 34.8k
HOPE (Sundermeyer et al., 2022) 5 options $\times$ $\times$ $\times$ 28 238 238 238
IPD (Kalra et al., 2024) 3 options 4 levels $\times$ $\times$ 20 30k 30k 30k
SenseShift6D (Ours) 5 levels 13 levels 9 levels 4 modes 5 166.4k 16.7k 665.9k
@@
Classic datasets. LineMOD (LM) (Hinterstoisser et al., 2012a) and its occlusion subset (LM-O) (Brachmann et al., 2014) pioneered RGB-D pose estimation benchmarks in controlled tabletop setups. Subsequent benchmarks such as T-LESS (Hodan et al., 2017), which provided a large collection of texture-less industrial objects, and YCB-V (Xiang et al., 2017), which offered cluttered scenes with temporal sequences, expanded the scope of evaluation. While these benchmarks provide robust evaluation regarding variations in objects and their poses, all were captured under fixed illumination and camera settings.
Initial exploration of environmental and sensor variation. TUD-L (Hodan et al., 2018b), HomebrewedDB (HB) (Kaskman et al., 2019), and HOPE (Sundermeyer et al., 2022) introduced variations in illumination conditions but retained fixed camera settings with auto-exposure enabled. The recent Industrial Parts Dataset (IPD) (Kalra et al., 2024) was the first benchmark to vary camera parameters, capturing four exposure levels under multiple lighting conditions. However, IPD focuses on texture-less industrial parts and does not explore variations in RGB gain or depth-sensor modes, thus limiting its suitability for adaptive-sensing research.
@@ 2.2 Adaptive Sensor Control Benchmark
Prior studies (Goudreault et al., 2023; Paul et al., 2022) have explored test-time adaptation of camera parameters to improve downstream vision performance, but these studies lacked explicit benchmarks in real-world. Recently, ImageNet-ES (Baek et al., 2024) and ImageNet-ES-Diverse (Baek et al., 2025) captured real-world scenes using a physical camera, systematically varying aperture, shutter speed, and ISO across multiple lighting conditions for image classification tasks. These benchmarks demonstrated that physical variation cannot be replicated by synthetic augmentations and dynamically selecting optimal camera parameters at test time can significantly increase the performance of lightweight models, enabling accuracy competitive with larger foundation models without additional training.
@@
Table 2: SenseShift6D Configurations for Data split.
[…]
Train-Def
50
Auto (1 RGB options,
1 depth option)
Auto
Auto
Default
RGB: 629, Depth: 629
1 RGB-D scene/pose
Train-Var
5, 25,
50, 75, 100
Auto + Manual
(41 RGB options,
1 depth options)
Auto, 2, 4, 19,
78, 312, 1250,
5000, 10000
Auto, 0, 32,
64, 96, 128
Default
RGB: 128.9k, Depth: 3.1k
225 RGB-D scenes/pose
Test-Def
50
Auto (1 RGB option,
1 depth option)
Auto
Auto
Default
RGB: 208, Depth: 208
1 RGB-D scenes/pose
Test-Var
5, 25,
50, 75, 100
Auto + Manual
(21 RGB options,
4 Depth options)
Auto, 9, 39,
156, 625, 2500
Auto, 16, 48,
80, 112
Default,
High Accuracy,
High Density,
Medium Density
RGB: 21.8k, Depth: 4.1k
480 RGB-D scenes/pose
@@ 3 SenseShift6D
Figure 3: Five household objects collected in SenseShift6D.
We introduce SenseShift6D, the first RGB-D benchmark that physically sweeps camera and lighting parameters, enabling rigorous evaluation of 6D object pose estimators under realistic environmental and sensor dynamics. Using an Intel RealSense D455 mounted in a darkroom, we systematically capture five representative household objects—spray, Pringles, tincase, sandwich, and mouse—selected to reflect diverse combinations of diverse RGB and geometric features (Figure 3), under auto-exposure, 13 RGB-exposure times, 9 RGB-gain settings, 4 depth-capture presets, and 5 illumination levels (details in Table 2). Our dataset consists of 166.4k RGB and 16.7k depth frames, which can yield 1,380 unique sensor-lighting permutations per object pose and 665.9k RGB-D combined scenes, the largest among the current benchamrks to our knowledge. Accurate ground truth is obtained via a rigidly mounted ChArUco board and careful hand-eye calibration.
@@
  • Train-Var: Expands Train-Def, covering (8 exposures $\times$ 5 gains + 5 auto-exposure) $\times$ 1 depth modes $\times$ 5 brightness levels (225 RGB-D scenes/pose).
@@
  • Test-Var: Expands Test-Def, featuring (5 exposures $\times$ 4 gains + 4 auto-exposure) configurations, spanning the same depth-mode and illumination range (480 RGB-D scenes/pose).
As all variations are physically captured rather than digitally simulated, the resulting images inherently retain authentic artifacts, such as non-linear color shifts, photon-shot noise, and depth-mode-specific errors, that synthetic augmentation methods cannot reproduce (Baek et al., 2024). SenseShift6D therefore offers the first realistic testbed for analyzing environment- and sensor-aware robustness, enabling side-by-side evaluation of per-frame adaptive-sensing policies against traditional data-centric training strategies.
@@ 3.1 Environmental and Sensor Variations
[…]
@@
Depth capture modes. The Intel RealSense D455 generates depth maps via stereo matching, estimating depth by aligning left and right camera images. This process relies on multiple internal parameters whose individual effects are complex and challenging to fine-tune directly. To manage this complexity, we leverage four predefined depth presets provided by Intel, each optimized for different usage scenarios (Hsu et al., 2023):
@@ 3.2 Dataset Collection
Table 3: GT annotation accuracy ($m$).
@@
For automatic collection, the turntable was remotely controlled, and deliberate delays were introduced after changing lighting brightness or sensor parameters, allowing sufficient stabilization time. We employed one smart IR remote-control hub to send commands to the turntable and another hub to control three smart bulbs. Both hubs received commands from a cloud-based server, which coordinated the entire data-capture sequence through an integrated control script (see Figure 4).
@@ 3.3 Data Processing
Ground-Truth annotation. For each pose, we used the 50%-brightness, averaged auto-exposure RGB frame and went through the annotation process below, resulting in about 2 mm errors.
  • Object masks: Initially, rough bounding boxes were generated using Grounding DINO (Liu et al., 2024), guided by shape-aware textual prompts tailored to each object. Boxes were manually reviewed and corrected if needed. The refined boxes were passed to FastSAM (Zhao et al., 2023) to generate precise instance-level masks.
  • Object poses: An initial camera-to-object transformation was provided by detecting the ChArUco markerboard rigidly attached to the turntable. The object poses were then refined precisely in 3D using the annotation utility (Gouda et al., 2022) in the BOP toolkit (Hodan & Sundermeyer, 2020) with depth maps captured in High Density mode. Finally, all refined poses were re-projected onto the RGB images and visually inspected to confirm alignment accuracy.
  • Accuracy of GT annotation: Following the protocol in (Hodan et al., 2017; Kaskman et al., 2019), we calculate the difference $\delta=d_{c}-d_{r}$, where $d_{c}$ denotes the captured depth map and $d_{r}$ denotes the rendered depth map using the GT pose. Differences over 5 cm are treated as outliers and omitted. Table 4 reports the mean $\mu_{\delta}$, standard deviation $\sigma_{\delta}$, and the mean and median of the absolute differences $\mu_{|\delta|}$ and $med_{|\delta|}$. High-accuracy depth sensing is employed for reliable evaluation of GT accuracy analysis.